OpenStudy (anonymous):

what the difference between .7 and .07

OpenStudy (anonymous):

Difference means to subtract

OpenStudy (anonymous):

just want to know if it is .7 is 70 and is .o7 is 7 cents

OpenStudy (anonymous):

unless they want you to say .7 = 7/100 and .07= 7/1000

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry I meant 7/10 and 7/100

OpenStudy (anonymous):

too many 0's lol! so sorry

OpenStudy (anonymous):

What context is the question in?

OpenStudy (anonymous):

in reference to money

OpenStudy (anonymous):

.7 is 70 cents and .07 is seven cents Assuming US dolalrs

sammixboo (sammixboo):

If you put 0.7 (0.70) in a money form if would be 70 cents, and if you put 0.07 in a money form it would be 7 cents...

OpenStudy (anonymous):

thank you all

sammixboo (sammixboo):

In fractions 0.7 = \[\frac{ 70 }{ 100 }\] and 0.07 = \[\frac{ 7 }{ 100 }\]

sammixboo (sammixboo):

You're welcome ;)